# Consistent Validation for Predictive Methods in Spatial Settings

### v2

@@ D.6.2 Rates of Convergence for Grid Prediction
Assumption D.15 (Grid prediction).
@@
Corollary D.16 (Bound on Estimation Error for Grid Prediction).
With the same assumptions as Theorem 5.1 and additionally Assumption D.15, with probability at least $1-\delta$
@@
See Section D.6.5 for a proof. The right-hand side of the bound is small as long as there is a validation point near each test point and the resolution of the map is high. If the available data for validation is generated i.i.d. as in Proposition 3.3, then right hand side becomes, up to logarithmic factors in $N^{\textup{{val}}}$, $\frac{1}{\sqrt{M^{\textup{{test}}}}}+(N^{\textup{{val}}})^{\min(-1/2,-1/d)}$
D.6.3 Statement and Discussion of Result for IID validation data and Grid Prediction
Corollary D.17 (Convergence of Grid Prediction with Independent and Identically Distributed Validation Data).
Suppose that $\mathcal{S}=[0,1]^{d}$, $S^{\textup{{val}}}_{n}\stackrel{{\scriptstyle\textup{{\tiny iid}}}}{{\sim}}P$ for $1\leq n\leq N^{\textup{{val}}}$ with $N^{\textup{{val}}}>1$ and $P$ has Lebesgue density lower bounded by $c>0$. Additionally, take the assumptions of Corollary D.16. Fix $\delta\in(0,1)$ and $k\in\{1,k^{\star}_{T_{2}}\}$. Then there exists a constant $K_{d,\delta,L,\Delta,c}$ that depends only on $d,\delta,L,c$ and $\Delta$ such that with probability at least $1-\delta$
@@ D.6.3 Statement and Discussion of Result for IID validation data and Grid Prediction
An upper bound on the bound in Corollary D.16 shows that for $k\in\{1,k^{\star}_{T_{2}}\}$ with probability $21-\delta/$ 2,
@@
Combining Equation 101 and Equation 102 via a union bound and using that $a+b\leq 2\max(a,b)$ completes the proof. ∎
This matches the bound proven in Portier et al. [2023, Proposition 3] which assumed that the test data was independent and identically distributed instead of on a regular grid. This rate of convergence is reasonably fast, particularly in low-dimensions. In particular ignoring the dependence of the bound on $M^{\textup{{test}}}$, which is reasonable as the number of test points in map prediction is often far larger than the number of available points for training and validation, it is faster than the minimax optimal rate of convergence for Lipschitz functions of $\theta((N^{\textup{{train}}})^{-1/(d+2)})$ discussed earlier. We therefore would expect both spatial nearest neighbors and $1$-nearest neighbors to perform well for model selection for grid prediction tasks if a fixed percentage of the data is used for training, and the remainder for validation. We emphasize we do not give a formal proof of this, just a heuristic argument suggesting why this should be the case. To give a formal proof would involve at least ensuring estimates of the risk estimation procedure hold uniformly over both sequences of predictive methods, and therefore involve additional assumptions.
D.6.4 General Prediction Tasks
For general $Q^{\textup{{test}}}$ we can combine Corollary D.14 together with Proposition 3.3 to get some sense of the rate of convergence of the spatial nearest neighbor method if the validation data is independent and identically distributed form a measure with density lower bounded on $[0,1]^{d}$ and the test task is fixed.
Corollary D.18 (Convergence of Spatial Nearest Neighbor with Independent and Identically Distributed Validation Data).
Suppose that $\mathcal{S}=[0,1]^{d}$, $S^{\textup{{val}}}_{n}\stackrel{{\scriptstyle\textup{{\tiny iid}}}}{{\sim}}P$ for $1\leq n\leq N^{\textup{{val}}}$ with $N^{\textup{{val}}}>1$ and $P$ has Lebesgue density lower bounded by $c>0$. Additionally, take the assumptions of Corollary D.14. Fix $\delta\in(0,1)$. Then there exists a constant $K_{d,\delta,L,\Delta,c}$ that depends only on $d,\delta,L,c$ and $\Delta$ such that with probability at least $1-\delta$
@@ D.6.4 General Prediction Tasks
Corollary D.14 implies that there exists a constant $K\geq 0$ such that with probability at least $21-\delta/$ 2
@@
Combining Equation 104 and Equation 105 completes the proof. ∎
In this case, again up to logarithmic factors, Corollary D.18 means that SNN converges at the optimal rate of convergence for Lipschitz functions. This means we do not necessarily expect to be able to distinguish between two sequences of predictive methods that converge at the minimax rate, but we might expect to distinguish between two sequences of predictive methods if one converges to the optimal predictor at much slower than the minimax rate. We again emphasize that we do not formally show this, and to do so would would involve at least ensuring estimates of the risk estimation procedure hold uniformly over both sequences of predictive methods, and therefore involve additional assumptions.
@@
D.6.5 Grid Prediction Proofs
In order to prove the claimed upper bound on grid prediction Corollary D.16 we use Theorem 5.1 together with an upper bound on the number of test points that lie within a ball of radius equal to the fill distance around any validation point. In order to do this, we will use that all the points in a grid are well-separated. We therefore begin by recalling the definition of a packing of a set, as well as a relationship between covering number and packing number.
Definition D.19 (Packing, Packing Number).
@@ D.6.5 Grid Prediction Proofs
Proposition D.20 (Packing and Covering Numbers Wainwright 2019, Lemma 5.5).
@@
We now restate and prove Corollary D.16. See D.16
@@
The second inequality follows from the first by Proposition D.12 and because $1\in T_{2}$. We therefore focus on proving the case $k=1$.
@@
The set $\{a/g:1\leq a\leq g\}^{d}$ is a $\frac{1}{g+\epsilon}$-packing for any $\epsilon>0$, and so by the first inequality in Proposition D.20 and Proposition D.3
@@
Applying Wainwright [2019, Lemma 5.7, Equation 5.9] and taking the limit as $\epsilon\to 0^{+}$
@@
Appendix E Additional Experimental Details
In this section, we provide additional details about the data, fitting procedures and validation procedures used in Section 6. Code used in experiments is available anonymously at: https://github.com/DavidRBurt/Consistent-Spatial-Validation. Code is almost all implemented in Python3 [Van Rossum and Drake, 2009] (with a small amount of r). Numpy is also heavily used for data generation and array manipulation [Harris et al., 2020].
In Section E.1 we give an overview of our method for estimating ground truth test risk in all experiments. In Section E.3 we describe details of the synthetic experiment described in Section 6.1. In Section E.4 we provide additional details on the air temperature data and tasks in Sections 6.2 and 6.5. In Section E.5 we provide additional details on the UK flat price prediction experiment described in Section 6.3, while in Section E.6 we provide additional details for the wind speed prediction task presented in Section 6.4.
E.1 Monte Carlo Estimation of Ground Truth Test Risk
@@
If we are willing to make the two assumptions above, we next show that we can reach (high probability) conclusions about the (true) relative quality of different estimators on a particular task if they pass a check: namely, we check if, for a small $\delta$ (e.g. $\delta=0.05$), the right-hand side of Equation 115 is smaller than twice the difference between how much closer the “good” estimator is to the estimate of ground truth than the “bad” estimate. To see why this check is sufficient, first observe the following two applications of the triangle inequality:
@@
Under the earlier assumptions, we see that Equation 120 is implied (with high probability) by
@@
E.2 Computational Considerations
E.2.1 Computational Complexity of our Method
We focus on the case $\mathcal{S}=[0,1]^{d}$ with nearest-neighbors implented using a $kd$-tree. Computation of the approximate fill distance is already discussed in Section D.5.4 and is shown to be $O(dN^{\textup{{val}}}\log N^{\textup{{val}}})$. Construction of a $kd$ tree on the validation data is also $O(dN^{\textup{{val}}}\log N^{\textup{{val}}})$. Once constructed, finding the nearest neighbor for each test point requires $O(\log N^{\textup{{val}}})$ computations of a $d$-dimensional Euclidean distance, meaning finding the neighbors is $O(dM^{\textup{{test}}}k\log N^{\textup{{val}}})$. Since $O(\log N^{\textup{{val}}})$ values of $k$ are tried in selecting $k$, this leads to a complexity not more than $O(dM^{\textup{{test}}}N^{\textup{{val}}}(\log N^{\textup{{val}}})^{2}).$ We expect further improvements could be made by storing the nearest neighbor set for intermediate values of $k$, but we do not pursue these.
E.2.2 Computational Setup Used
@@
E.2.3 Computational Cost of Synthetic Experiment
@@
E.2.4 Computational Cost of Bootstrapped Air Temperature Experiment
Data Generation
@@ E.2.4 Computational Cost of Bootstrapped Air Temperature Experiment
E.2.5 Computational Cost of UK House Price Experiment
Data Processing
@@ E.2.5 Computational Cost of UK House Price Experiment
E.2.6 Computational Cost of Wind Speed
Data Processing
@@ E.2.6 Computational Cost of Wind Speed
Running Experiment
Running the wind spped prediction task takes around 8.5 hours with 15 parallel jobs and has peak memory usage around 15GB.
E.2.7 Computational Cost of Real Data Air Temperature Experiment
Time to process the data has been previously described in Section E.2.4.
Running Experiment
@@ E.2.7 Computational Cost of Real Data Air Temperature Experiment
E.2.8 Computational Cost of Model Selection Experiment
@@
E.3 Risk Estimation on Synthetic Data
In this section, we provide additional details and figures for our risk estimation experiment on synthetic data presented in Section 6.1. Section E.3.1 describes the process by which we generate both datasets considered and shows an example of the covariate and response spatial fields for each problem (Figures 4 and 6). Section E.3.3 describes the procedure used to fit the predictive methods to each dataset. Section E.3.4 describes the implementation of the risk estimation procedures we compare. Section E.3.6 describes the metric reported, and Figures 5 and 7 show the (signed) relative error of each estimator as we vary the amount of validation data available.
E.3.1 Simulation Data Generating Process Details
@@
Figure 3: The validation sites (blue, clustered) for the first seed of the synthetic grid task. Panels from left to right and top to bottom represent $N^{\textup{{val}}}$ in the sequence $(250,500,1000,2000,4000,8000)$. Test sites (orange, gridded) are constant across panels.
@@
This is reminiscent of a Chinese restaurant process [Pitman, 2006, Section 3.1], but the weights are increased more slowly, leading to more clusters being formed and less large clusters typically.
@@
The first $1000$ points generated this way are taken to be the training data, and the remaining $N^{\textup{{val}}}$ points generated this way are the validation data. An example of the training and validation data generated through this process are shown in the top left of Figure 4.
@@
A small diagonal term (1e-12) is added to the diagonal of the covariance matrix to avoid numerical linear algebra errors. Sampling is performed using Tensorflow probability [Dillon et al., 2017]. We generate two covariates spatial processes via this process $\chi=(\chi^{(1)},\chi^{(2)})$.
@@
A small diagonal term (1e-12) is added to the diagonal of the covariance matrix to avoid numerical linear algebra errors. Sampling is performed using tensorflow probability [Dillon et al., 2017]. We generate two covariates via this process $X=(X^{(1)},X^{(2)})$.

### v3

@@ C.6.2 Rates of Convergence for Grid Prediction
Assumption C.16 (Grid prediction).
@@
Corollary C.17 (Bound on Estimation Error for Grid Prediction).
With the same assumptions as theorem 5.1 and additionally assumption C.16, with probability at least $1-\delta$
@@
See section C.6.5 for a proof. The right-hand side of the bound is small as long as there is a validation point near each test point and the resolution of the map is high. If the available data for validation is generated i.i.d. as in proposition 3.3, then right hand side becomes, up to logarithmic factors in $N^{\textup{{val}}}$, $\frac{1}{\sqrt{M^{\textup{{test}}}}}+(N^{\textup{{val}}})^{\min(-1/2,-1/d)}$
C.6.3 Statement and Discussion of Result for IID validation data and Grid Prediction
Corollary C.18 (Convergence of Grid Prediction with Independent and Identically Distributed Validation Data).
Suppose that $\mathcal{S}=[0,1]^{d}$, $S^{\textup{{val}}}_{n}\stackrel{{\scriptstyle\textup{{iid}}}}{{\sim}}P$ for $1\leq n\leq N^{\textup{{val}}}$ with $N^{\textup{{val}}}>1$ and $P$ has Lebesgue density lower bounded by $c>0$. Additionally, take the assumptions of corollary C.17. Fix $\delta\in(0,1)$ and $k\in\{1,k^{\star}_{T_{2}}\}$. Then there exists a constant $K_{d,\delta,L,\Delta,c}$ that depends only on $d,\delta,L,c$ and $\Delta$ such that with probability at least $1-\delta$
@@ C.6.3 Statement and Discussion of Result for IID validation data and Grid Prediction
An upper bound on the bound in corollary C.17 shows that for $k\in\{1,k^{\star}_{T_{2}}\}$ with probability $21-\delta/$ 2,
@@
Combining eq. 103 and eq. 104 via a union bound and using that $a+b\leq 2\max(a,b)$ completes the proof. ∎
This matches the bound proven in Portier et al. (2024, Proposition 4) which assumed that the test data was independent and identically distributed instead of on a regular grid. This rate of convergence is reasonably fast, particularly in low-dimensions. In particular ignoring the dependence of the bound on $M^{\textup{{test}}}$, which is reasonable as the number of test points in map prediction is often far larger than the number of available points for training and validation, it is faster than the minimax optimal rate of convergence for Lipschitz functions of $\theta((N^{\textup{{tr}}})^{-1/(d+2)})$ discussed earlier. We therefore would expect both spatial nearest neighbors and $1$-nearest neighbors to perform well for model selection for grid prediction tasks if a fixed percentage of the data is used for training, and the remainder for validation. We emphasize we do not give a formal proof of this, just a heuristic argument suggesting why this should be the case. To give a formal proof would involve at least ensuring estimates of the risk estimation procedure hold uniformly over both sequences of predictive methods, and therefore involve additional assumptions.
C.6.4 General Prediction Tasks
For general $Q^{\textup{{test}}}$ we can combine Corollary C.15 together with proposition 3.3 to get some sense of the rate of convergence of the spatial nearest neighbor method if the validation data is independent and identically distributed form a measure with density lower bounded on $[0,1]^{d}$ and the test task is fixed.
Corollary C.19 (Convergence of Spatial Nearest Neighbor with Independent and Identically Distributed Validation Data).
Suppose that $\mathcal{S}=[0,1]^{d}$, $S^{\textup{{val}}}_{n}\stackrel{{\scriptstyle\textup{{iid}}}}{{\sim}}P$ for $1\leq n\leq N^{\textup{{val}}}$ with $N^{\textup{{val}}}>1$ and $P$ has Lebesgue density lower bounded by $c>0$. Additionally, take the assumptions of corollary C.15. Fix $\delta\in(0,1)$. Then there exists a constant $K_{d,\delta,L,\Delta,c}$ that depends only on $d,\delta,L,c$ and $\Delta$ such that with probability at least $1-\delta$
@@ C.6.4 General Prediction Tasks
Corollary C.15 implies that there exists a constant $K\geq 0$ such that with probability at least $21-\delta/$ 2
@@
Combining eq. 106 and eq. 107 completes the proof. ∎
In this case, again up to logarithmic factors, corollary C.19 means that SNN converges at the optimal rate of convergence for Lipschitz functions. This means we do not necessarily expect to be able to distinguish between two sequences of predictive methods that converge at the minimax rate, but we might expect to distinguish between two sequences of predictive methods if one converges to the optimal predictor at much slower than the minimax rate. We again emphasize that we do not formally show this, and to do so would would involve at least ensuring estimates of the risk estimation procedure hold uniformly over both sequences of predictive methods, and therefore involve additional assumptions.
@@
C.6.5 Grid Prediction Proofs
In order to prove the claimed upper bound on grid prediction corollary C.17 we use theorem 5.1 together with an upper bound on the number of test points that lie within a ball of radius equal to the fill distance around any validation point. In order to do this, we will use that all the points in a grid are well-separated. We therefore begin by recalling the definition of a packing of a set, as well as a relationship between covering number and packing number.
Definition C.20 (Packing, Packing Number).
@@ C.6.5 Grid Prediction Proofs
Proposition C.21 (Packing and Covering Numbers Wainwright 2019, Lemma 5.5).
@@
We now restate and prove corollary C.17. See C.17
@@
The second inequality follows from the first by proposition C.13 and because $1\in T_{2}$. We therefore focus on proving the case $k=1$.
@@
The set $\{a/g:1\leq a\leq g\}^{d}$ is a $\frac{1}{g+\epsilon}$-packing for any $\epsilon>0$, and so by the first inequality in proposition C.21 and proposition C.3
@@
Applying Wainwright (2019, Lemma 5.7, Equation 5.9) and taking the limit as $\epsilon\to 0^{+}$
@@
Appendix D ADDITIONAL EXPERIMENTAL DETAILS
In this section, we provide additional details about the data, fitting procedures and validation procedures used in section 6. Code used in experiments is available anonymously at: https://github.com/DavidRBurt/Consistent-Spatial-Validation. Code is almost all implemented in Python3 (Van Rossum and Drake, 2009) (with a small amount of r). Numpy is also heavily used for data generation and array manipulation (Harris et al., 2020).
In section D.1 we give an overview of our method for estimating ground truth test risk in all experiments. In section D.3 we describe details of the synthetic experiment described in section 6.1. In section D.4 we provide additional details on the air temperature data and tasks in sections 6.2 and 6.5. In section D.5 we provide additional details on the UK flat price prediction experiment described in section 6.3, while in section D.6 we provide additional details for the wind speed prediction task presented in section 6.4.
D.1 Monte Carlo Estimation of Ground Truth Test Risk
@@
If we are willing to make the two assumptions above, we next show that we can reach (high probability) conclusions about the (true) relative quality of different estimators on a particular task if they pass a check: namely, we check if, for a small $\delta$ (e.g. $\delta=0.05$), the right-hand side of eq. 117 is smaller than twice the difference between how much closer the “good” estimator is to the estimate of ground truth than the “bad” estimate. To see why this check is sufficient, first observe the following two applications of the triangle inequality:
@@
Under the earlier assumptions, we see that eq. 122 is implied (with high probability) by
@@
D.2 Computational Considerations
D.2.1 Computational Complexity of our Method
We focus on the case $\mathcal{S}=[0,1]^{d}$ with nearest-neighbors implemented using a $k$-d tree.
Proposition D.1 (Computational Complexity of Spatial Nearest Neighbors).
Suppose $\mathcal{S}=[0,1]^{d}$ then the spatial nearest-neighbor estimator can be computed in $O(dM^{\textup{{test}}}N^{\textup{{val}}}+M^{\textup{{test}}}N^{\textup{{val}}}\log N^{\textup{{val}}})$ algebraic operations.
Proof.
We have already shown that computation of the approximate fill distance is $O(dN^{\textup{{val}}}\log N^{\textup{{val}}})$ to be in section C.5.4. While we use $k$-d trees to implement nearest neighbors in practice, we consider the complexity of a direct implementation of nearest neighbors. In this case, computing all pairwise distances between test and validation points has complexity $O(dM^{\textup{{test}}}N^{\textup{{val}}})$. Once these distances are computed, the distances for each validation point can be sorted, and the original indices can then be used to recover the nearest neighbors. The sorting operation has complexity $O(M^{\textup{{test}}}N^{\textup{{val}}}\log N^{\textup{{val}}})$. Given the sorted list of distances and corresponding indices computing the weights needed for computing the bound for each $k$ points involves counting the number of times each training index appears in a sub-array consisting of the first $k$ columns. This can be done by scanning the array from left to right and keeping track of partial sums in $O(N^{\textup{{val}}}M^{\textup{{test}}})$. And given the weights, we can compute the bound for all of the $\log N^{\textup{{val}}}$ many values of $k$ each in time $N^{\textup{{val}}}$, giving a complexity of not more than $O(N^{\textup{{val}}}\log N^{\textup{{val}}})$. Taking all of these steps together, we see the overall complexity is, $O(dM^{\textup{{test}}}N^{\textup{{val}}}+M^{\textup{{test}}}N^{\textup{{val}}}\log N^{\textup{{val}}}).$
Figure 5: Computational time (in seconds) for computation of the three estimates of validation error in the synthetic experiment with data at a single test point (left) and on a regular grid (right). The mean is shown with the main line, while the maximum and minimum are indicated by the vertical bars. For both datasets, our method takes significantly longer to compute than the baselines But even with the largest number of validation points used, our method takes under a minute to compute.
D.2.2 Computational Setup Used
@@
D.2.3 Computational Cost of Synthetic Experiment
@@
Computational Cost of Validation Methods.
Figure 5 shows the time taken to compute each estimator of the validation risk as a function of the number of validation points. Our method takes longer than baselines, but is still fast to compute in practice.
D.2.4 Computational Cost of Bootstrapped Air Temperature Experiment
Data Generation.
@@ D.2.4 Computational Cost of Bootstrapped Air Temperature Experiment
Computational Cost of Validation Methods.
Metro Prediction: For geographically weighted regression, the holdout took a mean time of 8.66 seconds with standard deviation of 3.93 seconds. 1NN took a mean time of 7.98 with standard deviation 3.53 seconds. SNN took a mean time of 29.89 seconds with a standard deviation of 11.48 seconds. For kernel ridge regression the holdout took a mean time of 3.06 seconds with standard deviation 0.93 seconds. 1NN took a mean time of 3.35 seconds with a standard deviation of 0.94 seconds. SNN took a mean time of 19.87 seconds and standard deviation 3.73 seconds.
Grid Prediction: For geographically weighted regression, the holdout took a mean time of 3.88 seconds with standard deviation of 0.60 seconds. 1NN took a mean time of 5.83 with standard deviation 0.59 seconds. SNN took a mean time of 171.49 seconds with a standard deviation of 2.75 seconds. For kernel ridge regression the holdout took a mean time of 0.72 seconds with standard deviation 0.02. 1NN took a mean time of 2.57 seconds and standard deviation of 0.026 seconds. SNN took a mean time of 169.91 seconds and standard deviation 3.77 seconds.
D.2.5 Computational Cost of UK House Price Experiment
Data Processing.
@@ D.2.5 Computational Cost of UK House Price Experiment
Computational Cost of Validation Methods.
The holdout took a mean time of 44.97 seconds with standard deviation 5.77. 1NN took a mean time of 43.37 seconds and standard deviation of 5.02 seconds. SNN took a mean time of 66.18 seconds and standard deviation 8.49 seconds.
D.2.6 Computational Cost of Wind Speed
Data Processing.
@@ D.2.6 Computational Cost of Wind Speed
Running Experiment.
Running the wind speed prediction task takes around 8.5 hours with 15 parallel jobs and has peak memory usage around 15GB.
Computational Cost of Validation Methods.
The holdout took a mean time of 0.14 seconds with standard deviation 0.09. 1NN took a mean time of 0.50 seconds and standard deviation of 0.33 seconds. SNN took a mean time of 51.71 seconds and standard deviation 15.71 seconds.
D.2.7 Computational Cost of Real Data Air Temperature Experiment
Time to process the data has been previously described in section D.2.4.
Running Experiment.
@@ D.2.7 Computational Cost of Real Data Air Temperature Experiment
Computational Cost of Validation Methods.
Metro Prediction: For geographically weighted regression, the holdout took 2.31 seconds. 1NN took 2.31. SNN took 2.32 seconds. For kernel ridge regression the holdout took 1.75 seconds. 1NN took 1.75 seconds. SNN took 1.75 seconds.
Grid Prediction: For geographically weighted regression, the holdout took 2.30 seconds. 1NN took 3.54. SNN took 83.1 seconds. For kernel ridge regression the holdout took 1.70 seconds. 1NN took 3.00 seconds. SNN took 82.2 seconds.
D.2.8 Computational Cost of Model Selection Experiment
@@
D.3 Risk Estimation on Synthetic Data
In this section, we provide additional details and figures for our risk estimation experiment on synthetic data presented in section 6.1. Section D.3.1 describes the process by which we generate both datasets considered and shows an example of the covariate and response spatial fields for each problem (figs. 7 and 9). Section D.3.3 describes the procedure used to fit the predictive methods to each dataset. Section D.3.4 describes the implementation of the risk estimation procedures we compare. Section D.3.6 describes the metric reported, and figs. 8 and 10 show the (signed) relative error of each estimator as we vary the amount of validation data available.
D.3.1 Simulation Data Generating Process Details
@@
Figure 6: The validation sites (orange diamonds, clustered) for the first seed of the synthetic grid task. Panels from left to right and top to bottom represent $N^{\textup{{val}}}$ in the sequence $(250,500,1000,2000,4000,8000)$. Test sites (blue triangles, gridded) are constant across panels.
@@
This is reminiscent of a Chinese restaurant process (Pitman, 2006, Section 3.1), but the weights are increased more slowly, leading to more clusters being formed and less large clusters typically.
@@
The first $1000$ points generated this way are taken to be the training data, and the remaining $N^{\textup{{val}}}$ points generated this way are the validation data. An example of the training and validation data generated through this process are shown in the top left of fig. 7.
@@
A small diagonal term (1e-12) is added to the diagonal of the covariance matrix to avoid numerical linear algebra errors. Sampling is performed using Tensorflow probability (Dillon et al., 2017). We generate two covariates spatial processes via this process $\chi=(\chi^{(1)},\chi^{(2)})$.
@@
A small diagonal term (1e-12) is added to the diagonal of the covariance matrix to avoid numerical linear algebra errors. Sampling is performed using tensorflow probability (Dillon et al., 2017). We generate two covariates via this process $X=(X^{(1)},X^{(2)})$.